OpenStudy (anonymous):

Which of the following equations is of a parabola with a vertex at (3, 0)? I think it is y = x^2+ 3

OpenStudy (anonymous):

no it wouldn't be that because sense 3 is an x when you graph y = x^2 + 3 it would go up the y-axis not over on the x-axis so making a equation like on the last problem you did do you think you could figure it out?

OpenStudy (anonymous):

it would be an equation like y = (x+1)^2+5 with the correct numbers of course

OpenStudy (anonymous):

y = x^2- 3

OpenStudy (anonymous):

y = (x + 3)^2 This one srry

OpenStudy (anonymous):

very close but remember when you take that form and change it into a set of points on the graph you have to switch the sign in front of the first one

OpenStudy (anonymous):

y = (x - 3)^2

OpenStudy (anonymous):

YES!!! Very good.

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

You are very welcome.